OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5).

OpenStudy (chriss):

y=mx+b \[m=\frac{5-(-4)}{2-(-1)}\] \[m=\frac{9}\] m=3 5=4(2)+b 5=8+b -3=b y=3x-3

OpenStudy (precal):

y=3x-1

OpenStudy (anonymous):

 use the formula. .